Sol Summary – April 16th

Crew 297 Sol Summary Report 16-04-2024

Sol: 02

Summary Title: Aux Power has Been Recovered!

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

Janus I woke up today with high spirits and a great sense of adventure – a necessary state of mind to be in for the demanding EVA prepared today to recover a lost [inert] NPS. Last night an auxiliary power source in the form of an NPS finally landed on the Red Planet. However, due to strong winds the NPS did not arrive at its expected landing location. The location was narrowed down to a radius of 300 m within the Hab and crew members Matt S., Matt L., and Sean went through heroic efforts to try to locate it. After hours of unsuccessful efforts, remote sensing from Martian satellites were able to locate the NPS, and a more precise location was uncovered. The crew quickly located the device and continued the remainder of their EVA plan nominally.

In spite of the unexpectedly taxing NPS recovery efforts, the crew remains optimistic about the days ahead. After report writing, the crew now winds down, ready for any challenges Sol 3 may bring.

Look Ahead Plan:

An EVA is planned tomorrow morning to replicate the [inert] NPS search and recovery with a different EVA team. After a (hopefully prompt) recovery effort the EVA will continue with Sarah commencing her gamma-ray portable spectrometer research. The rest of the crew will recover from their valiant efforts today and continue their nominal duties including preparing and conducting IVA research.

Anomalies in work:

Suit 3 (Loose XLR 3 pin connector): fixed and functionally tested today (via EVA).
Suit 5 (ventilation electrical connection): investigated and likely fixed. To be further examined/tested in the coming days.
Morning (data from 0800): calm winds (4 mph, gusting 11 mph), chilly (50F), and clear skies (0% cloud coverage)
Noon: moderate winds (7 mph, gusting 15 mph), warm (62F), and clear skies (0% cloud coverage)
Afternoon (1700): moderate winds (5 mph, gusting 17 mph), warm (70F), and clear skies (0% cloud coverage)
Crew Physical Status: Crew members that participated in the 4-hour EVA are physically drained. However, overall the crew is in good spirits and health.

EVA: Although strenuous, the only EVA today was a success in that the NPS was found (and rehidden for tomorrow’s EVA) and the crew returned safely.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (1 EVA)
EVA request (1 EVA)
Support Requested: None requested.

Copyright © The Mars Society. All rights reserved. | Main Site